Find the mean and standard deviation for each binomial random variable:

a. n = 60, π = .65 (Round your mean value to 2 decimal places and standard deviation to 4 decimal places.)


Answer by Boreal(15235)

mean is nπ=39.00
variance is nπ(1-π)=39*0.35=13.65
sd is sqrt(V)=3.6946